OpenStudy (anonymous):

Determine whether F is the gradient of some function f. If it is, find such a function f. 1. F(x,y)=(e^y)i+((xe^y)+y)j 2. F(x,y)=(2xyz)i+(x^2*z)j+(x^2*y+1)k

OpenStudy (anonymous):

if its a gradient , it should satisfy: \[\frac{\partial e^y}{\partial y}=\frac{\partial (xe^y +y)}{\partial x}\] which they do

OpenStudy (anonymous):

Gotcha. So what are the steps/formula for finding the function?

OpenStudy (anonymous):

Let's say its gradient of some function G

OpenStudy (anonymous):

Then \[\frac{\partial G}{\partial x}=e^y\]

OpenStudy (anonymous):

Integrating respect to x: \[G=xe^y +g(y)\] differentiating this respect to y: \[\frac{\partial G}{\partial y} =xe^y+g'(y)\]

OpenStudy (anonymous):

but dG/dy= xe^y)+y so \[xe^y+g'(y)=xe^y+y\]

OpenStudy (anonymous):

from here you see that g'(y)=y

OpenStudy (anonymous):

so integrating: g(y) = 1/2y^2 so G(x,y)=xe^y+1/2y^2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Hmmm, can you write out your steps in words? I understand to a degree. I just wanna write this out on paper so I can recheck the steps.

OpenStudy (anonymous):

I did write it in words

OpenStudy (anonymous):

you can write this on paper like this and you will be fine

OpenStudy (anonymous):

more clear impossible

OpenStudy (anonymous):

Alright I think I got it then, thanks.

OpenStudy (anonymous):

yw